Shift report 20020108 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 08 2002 - 23:18:00 EST

  • Next message: Apache: "Shift report 20020108 16:00-24:00"

    ------------------------------------------------------------
    Date: Tue Jan  8 23:18:00 2002, Shift: 16:00-24:00
    Supervisor: IGB, with: mads
    Log book#12 pages 29 - 30
    
    Shift Summary:
    ------------------------------------------------------------
    Machine Status:
    Had one store from 1730 until 2130.  The beam quality was not 
    terribly good, but as it was impossible to get a new fill, decided that 
    not good beam was still better than no beam.
    No beam since 2130, but the machine is being filled for a ramp.
    
    Achievements:
    Finished 100k events with the inner innelastic counters out 
    of the trigger.
    Replaced these signals in the discriminator 
    
    Detector Problems:
    
    Flemming fixed a small timing problem in the trigger, 
    otherwise no problems were noticed.
    
    We were unable to submit the shift check using the Web form.  
    We decided that it was not worth interupting Flemmings 
    (late, because of the timing problem) supper for 
    this minor problem. 
    
    Had some problems getting the DAQ to work properly.  
    FV fixed this, though I'm not completely sure how, exactly, he managed 
    this.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 08 2002 - 23:18:54 EST